OpenStudy (anonymous):

What is the sum of the series below? 6 ∑(7i-2) i=1

OpenStudy (anonymous):

135

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

evaluate for each value of i and add

OpenStudy (anonymous):

(7-2)+(14-2)+(21-2)+(28-2)+(35-2)+(42-2)

OpenStudy (australopithecus):

6 ∑(7i-2) = (7(1) - 2) + (7(2) - 2) + (7(3) - 2) + (7(4) - 2)+ (7(5) - 2) + (7(6) - 2) i = 1

OpenStudy (australopithecus):

there is a formula you can use to calculate this but don't expect to have to do this very much so the formulas are more of a novelty usually at least they were for me